# Random pairing MLE for estimation of item parameters in Rasch model

### v1

@@ Abstract
The Rasch model, a classical model in the item response theory, is widely used in psychometrics to model the relationship between individuals’ latent traits and their binary responses on assessments or questionnaires. In this paper, we introduce a new likelihood-based estimator—random pairing maximum likelihood estimator ($\mathsf{RP\text{-}MLE}$) and its bootstrapped variant multiple random pairing MLE ($\mathsf{MRP\text{-}MLE}$) that faithfully estimate the item parameters in the Rasch model. The new estimators have several appealing features compared to existing ones. First, both work for sparse observations, an increasingly important scenario in the big data era. Second, both estimators are provably minimax optimal in terms of finite sample $\ell_{\infty}$ estimation error. Lastly, $\mathsf{RP\text{-}MLE}$ admits precise distributional characterization that allows uncertainty quantification on the item parameters, e.g., construction of confidence intervals of the item parameters. The main idea underlying $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ is to randomly pair user-item responses to form item-item comparisons. This is carefully designed to reduce the problem size while retaining statistical independence. We also provide empirical evidence of the efficacy of the two new estimators using both simulated and real data.
@@ 1 Introduction
The item response theory (IRT) [ER13] is a framework widely used in psychometrics to model the relationship between individuals’ latent traits (such as ability or personality) and their responses on assessments or questionnaires. It is particularly useful in the development, analysis, and scoring of tests and assessments; see a recent survey [CLLY21] for a statistical account of IRT.
@@
Among statistical models in IRT, the Rasch model [Ras60] is a simple but fundamental one for modeling binary responses. Specifically, for a user $t$ (e.g., test-taker) and an item $i$ (e.g., test problem), Rasch model assumes that the response user $t$ has to item $i$ is binary and obeys
@@
$\mathbb{P}\left[\text{user $t$ ``beats'' item $i$}\right]=\frac{e^{\zeta_{t}^{\star}}}{e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}}},$
@@
where $\zeta_{t}^{\star},\theta_{i}^{\star}\in\mathbb{R}$ are latent traits of user $t$ and item $i$, respectively. The term “beats” here refers to positive responses, such as answering an exam question correctly, writing a favorable review of a product, etc. We also call this response a comparison between user $t$ and item $i$.
@@
In this paper, we focus on estimating the item parameters $\bm{\theta}^{\star}$, which is one of the four main statistical tasks surrounding the Rasch model (or IRT more generally) listed in [CLLY21]. Estimating the item parameters is also quite useful in practice. For instance, in education testing, $\bm{\theta}^{\star}$ could reveal the difficulty of the exam questions, while in product reviews, $\bm{\theta}^{\star}$ could reveal the popularity of the products. Various methods have been proposed for estimating the item parameters $\bm{\theta}^{\star}$, including the joint maximum likelihood estimator (JMLE), the marginal maximum likelihood estimator, the conditional maximum likelihood estimator (CMLE), as well as the spectral estimator recently proposed in [NZ22, NZ23]. We refer readers to a recent article [Rob21] for comparisons between different item parameter estimation methods. However, three main gaps remain in tackling item estimation in the Rasch model:
@@
  • Non-asymptotic guarantee. Apart from the recently proposed spectral estimator [NZ22, NZ23], most theoretical guarantees for the likelihood-based estimators are asymptotic. Since all the estimation procedures are necessarily applied with finite samples, the asymptotic guarantee alone fails to inform practitioners about the performance of different estimators when working with a limited number of samples.
@@
  • Sparse observations. It is not uncommon to encounter situations where each user only responds to a handful of questions (or items). This brings the challenge of incomplete and sparse observations. Among the likelihood based estimators, CMLE is known to have numerical issues with incomplete observations [Mol95]. While the spectral estimator [NZ22, NZ23] is capable of handling incomplete observations, it still requires the observations to be relatively dense. We will elaborate on this point later.
@@
  • Uncertainty quantification. Beyond estimation, uncertainty quantification on the item parameters is central to realizing the full potential of the Rasch model. However, existing results do not address this problem under sparse observations. An exception is the recent work [CLOX23], which is based on joint estimation and inference on the item parameters $\bm{\theta}^{\star}$ and the user parameters $\bm{\zeta}^{\star}$. Their sampling scheme is restrictive and requires a relatively dense sampling rate.
In light of the above gaps, we raise the following natural question:
@@ 1.1 Main contributions
The main contribution of our work is the proposal of a novel estimator named random pairing maximum likelihood estimator ($\mathsf{RP\text{-}MLE}$ in short) that achieves the two desiderata listed above.
In essence, $\mathsf{RP\text{-}MLE}$ compiles user-item comparisons to item-item comparisons by randomly pairing responses of the same user to different items. This pairing procedure is carefully designed to extract information of the item parameters while retaining statistical independence. After this compilation step, item parameters $\bm{\theta}^{\star}$ are estimated by the MLE $\widehat{\bm{\theta}}$ given the item-item comparisons.
Even when the observations are extremely sparse, $\mathsf{RP\text{-}MLE}$ achieves the following:
  • Regarding estimation, we show that both $\mathsf{RP\text{-}MLE}$ and its bootstrapped version enjoy optimal finite sample $\ell_{\infty}$ error guarantee. Compared to the conventional $\ell_{2}$ error guarantee, the $\ell_{\infty}$ guarantee, as an entrywise guarantee is more fine-grained. Consequently, we also show that $\mathsf{RP\text{-}MLE}$ can recover the top-$K$ items with minimal sample complexity.
  • While the optimal $\ell_{\infty}$ error guarantee directly yields optimal $\ell_{2}$ guarantee, such guarantee is only correct in an order-wise sense. We provide a refined finite-sample $\ell_{2}$ error guarantee of $\mathsf{RP\text{-}MLE}$ that is precise even in the leading constant.
  • Supplementing the estimation guarantee, we also build an inferential framework based on $\mathsf{RP\text{-}MLE}$ $\widehat{\bm{\theta}}$. More specifically, we precisely characterize the asymptotic distribution of $\widehat{\bm{\theta}}$. This result facilitates several inferential tasks such as hypothesis testing and construction of confidence regions of $\bm{\theta}^{\star}$.
@@ 1.2 Prior art
The item response theory is a popular statistical framework for modeling response data. It often involves a probabilistic model that links categorical responses to latent traits of both users and items. Early endeavors include [Ras60] that introduces the Rasch model studied herein and [LNB68] that describes a more general framework using parametric models. Popular IRT models include the Rasch model, the two-parameter model (2PL), and the three-parameter logistic model (3PL). As response data widely appears in real life, IRT finds application in numerous fields including educational assessment [DC10], psychometrics [LNB68], political science [VHSA20], and medical assessment [FBC05]. See also [CLLY21] for an overview of IRT from statisticians’ perspective.
@@
An important statistical question in the Rasch model is to estimate the latent parameters of the items. As the Rasch model is an explicit probabilistic model, many methods are based on the principle of maximum likelihood estimation. For instance, marginal MLE assumes a prior on the user parameters that is either given or optimized within a parametric distribution family. The item parameter is then estimated by maximizing the marginal likelihood. A drawback is that MMLE relies on a good prior. On the other hand, joint MLE (JMLE) makes no distributional assumption and maximizes the joint likelihood w.r.t. both the item and user parameters. However, it is not consistent for estimating the item parameters when the number $m$ of items is fixed [Gho95]. Interested readers may also consult [Lin99] for an overview of other classical estimators.
Several methods are more relevant to our proposed estimator $\mathsf{RP\text{-}MLE}$ as they follow a similar philosophy to form item-item pairs from user-item responses. Pseudo MLE (PMLE) [Zwi95] maximizes the sum of the log-likelihood of all pairs of responses from the same users to different items. However, due to the dependency issue, no satisfying finite sample performance guarantee has been established. Another related approach is the spectral method, in which a Markov chain on the space of items is formed and the item parameters are estimated via the stationary distribution of the Markov chain. The most recent works in this category are [NZ22, NZ23], which essentially use the same idea as pseudo MLE in forming item-item comparisons.
@@
An informed reader may realize that the Rasch model resembles the Bradley-Terry-Luce (BTL) model [Luc59, BT52] in the ranking literature. Indeed, one can view the Rasch model as a special case of the BTL model that distinguishes the two groups of users and items, and only allows inter-group comparisons. There has been a recent surge in interest in studying top-$K$ ranking in the BTL model [CS15, JKSO16, CFMW19, CGZ22, GSZ23] and its extensions [FHY22, FLWY22, FLWY23], especially under sparse observations of the pairwise responses. Most notably, under a uniform sampling scheme, [CFMW19] shows that (regularized) MLE and spectral methods are both optimal in top-$K$ ranking and [GSZ23] provides inference results for both methods. Going beyond uniform sampling, [Che23, LSR22] investigate the performance of MLE in the BTL model with a general comparison graph and later [YCOM24] improves the analysis to show the optimality of MLE for the BTL model in both uniform and semi-random sampling.
@@
For a positive integer $n$, we denote $[n]=\{1,2,\ldots,n\}$. For any $a,b\in\mathbb{R}$, $a\wedge b$ means the minimum of $a,b$ and $a\vee b$ means the maximum of $a,b$. We use $\overset{\mathrm{d}}{\rightarrow}$ to denote convergence in distribution. For a symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$, we use $\lambda_{1}(\bm{A})\geq\lambda_{2}(\bm{A})\geq\cdots\geq\lambda_{n}(\bm{A})$ to denote its eigenvalues and $\bm{A}^{\dagger}$ to denote its Moore-Penrose pseudo-inverse. For symmetric matrices $\bm{A},\bm{B}\in\mathbb{R}^{n\times n}$, $\bm{A}\preceq\bm{B}$ means $\bm{B}-\bm{A}$ is positive semidefinite, i.e., $\bm{v}^{\top}(\bm{B}-\bm{A})\bm{v}\geq 0$ for any $\bm{v}\in\mathbb{R}^{n}$. We use $\bm{e}_{i}$ to denote the standard unit vector with $1$ at $i$-th coordinate and 0 elsewhere. Unless specified otherwise, $\log(\cdot)$ denotes the natural log.
2 Problem setup and new estimator $\mathsf{RP\text{-}MLE}$
In this section, we first introduce the formal setup of the item parameter estimation problem in the Rasch model. Then we present the newly proposed estimator $\mathsf{RP\text{-}MLE}$ along with the rationale behind its development.
@@ 2.1 Problem setup
To model sparse observations/comparisons, we assume that $\mathbb{P}[(t,i)\text{ is compared}]=p$ independently for every $(t,i)\in[n]\times[m]$. To put it in the language of graph theory, we denote the associated bipartite comparison graph to be $\mathcal{G}_{X}=(\mathcal{V}_{X},\mathcal{E}_{X})$, where $\mathcal{V}_{X}$ consists of $n$ users and $m$ items. Then essentially, we are assuming that the bipartite graph follows an Erdős-Rényi random model. [^1] Alternatively, we can assume each user responds to $mp$ items uniformly at random. Our estimator and performance guarantee continue to work in this sampling scheme.
Before moving on, we introduce the condition numbers that characterize the range of the latent traits. Let $\kappa_{1}$, $\kappa_{2}$, and $\kappa$ be defined by $\log\;(\kappa_{1})=\max_{ij}\{|\theta_{i}^{\star}-\theta_{j}^{\star}|\}$, $\log(\kappa_{2})=\max_{ti}\{|\zeta_{t}^{\star}-\theta_{i}^{\star}|\}$, and $\kappa=\max\{\kappa_{1},\kappa_{2}\}$, respectively.
@@ 2.2 Random pairing maximum likelihood estimator
Algorithm 1 Random Pairing Maximum Likelihood Estimator ($\mathsf{RP\text{-}MLE}$)
@@
      Randomly split the $n_{t}$ problems taken by tester $t$ into $\lfloor n_{t}/2\rfloor$ pairs of problems.
@@
    Let $\mathcal{E}_{Y}$ be a set of edges defined by $\mathcal{E}_{Y}\coloneqq\{(i,j):\sum_{t=1}L_{ij}^{t}\geq 1\}$ and let $\mathcal{G}_{Y}=([m],\mathcal{E}_{Y})$. For each $(i,j)\in\mathcal{E}_{Y}$, let $L_{ij}\coloneqq\sum_{t=1}^{n}L_{ij}^{t}$ and $Y_{ij}\coloneqq(1/L_{ij})\sum_{\{t:L_{ij}^{t}=1\}}Y_{ij}^{t}$.
@@
    Compute MLE on $Y_{ij}$, i.e.,
    $\widehat{\bm{\theta}}\coloneqq\arg\min_{\bm{\theta}:\bm{1}_{m}^{\top}\bm{\theta}=0}\mathcal{L}(\bm{\theta})$
    where
@@
In this section, we present our main method $\mathsf{RP\text{-}MLE}$; see Algorithm 1. The algorithm can be divided into two parts. The first part—Steps 1 and 2—uses random pairing to compile the observed user-item responses $\bm{X}\in\mathbb{R}^{n\times m}$ to item-item comparisons $\bm{Y}\in\mathbb{R}^{m\times m}$. The second part—Steps 3 and 4—computes a standard MLE on the item-item comparisons. Some intuitions regarding the development of $\mathsf{RP\text{-}MLE}$ are in order.
@@
The idea of pairing is that by matching the responses $X_{ti}$ with $X_{tj}$, we form a comparison between items $i$ and $j$ to directly extract information of item parameters $\theta_{i}^{\star}$ and $\theta_{j}^{\star}$. More specifically, the item-item comparisons $\bm{Y}$ follow the Bradley-Terry-Luce model [BT52, Luc59],i.e., $\mathbb{P}[Y_{ij}^{t}]=e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})$; see Section 4.1 for a formal argument.
@@
More importantly, the pairing is performed in a disjoint fashion. This ensures that all constructed item-item comparisons $Y_{ij}^{t}$ are independent with each other; see Section 4.1 for a formal statement. This is the key ingredient that enables us to improve over previous implementation of item-item comparisons, such as pseudo-likelihood [Cho82, Zwi95] and spectral methods [NZ22, NZ23].
@@
A drawback of this random pairing is that it potentially induces a loss of information since not every possible pairing is considered. Once $X_{ti}$ is paired with $X_{tj}$, we cannot pair $X_{ti}$ with another response $X_{tl}$. That being said, we will later show that the $\ell_{\infty}$ error of $\mathsf{RP\text{-}MLE}$ is still rate-optimal up to logarithmic factors. Hence the loss of information can at most incur a small constant factor in terms of estimation error. Nevertheless, we provide a remedy to this phenomenon in $\mathsf{MRP\text{-}MLE}$ (Algorithm 2) by running (in other words, bootstrapping) the $\mathsf{RP\text{-}MLE}$ multiple times with different random data splitting and averaging the resulting estimates. $\mathsf{MRP\text{-}MLE}$ trivially enjoys the same theoretical guarantee (cf. Theorem 1) while improving the estimation error in practice over $\mathsf{RP\text{-}MLE}$. See Figure 4 in Section 5.1 for the empirical evidence.
Algorithm 2 Multiple Random Pairing Maximum Likelihood Estimator ($\mathsf{MRP\text{-}MLE}$)
@@
    Let $n_{\mathrm{split}}$ be the number of runs. For $i=1,\ldots,n_{\mathrm{split}}$, run $\mathsf{RP\text{-}MLE}$ (Algorithm 1), each time with an independent random splitting in Step 1. Let the estimated latent scores be $\widehat{\bm{\theta}}^{(i)}$.
@@
    $\widehat{\bm{\theta}}=\frac{1}{n_{\mathrm{split}}}\sum_{i=1}^{n_{\mathrm{split}}}\widehat{\bm{\theta}}^{(i)}.$
@@
    Estimate the top-$K$ items by selecting the top-$K$ entries of $\widehat{\bm{\theta}}$.

### v2

@@ Abstract
The Rasch model, a classical model in the item response theory, is widely used in psychometrics to model the relationship between individuals’ latent traits and their binary responses to assessments or questionnaires. In this paper, we introduce a new likelihood-based estimator—random pairing maximum likelihood estimator ($\mathrm{RP\text{-}MLE}$) and its bootstrapped variant multiple random pairing MLE ($\mathrm{MRP\text{-}MLE}$) which faithfully estimate the item parameters in the Rasch model. The new estimators have several appealing features compared to existing ones. First, both work for sparse observations, an increasingly important scenario in the big data era. Second, both estimators are provably minimax optimal in terms of finite sample $\ell_{\infty}$ estimation error. Lastly, both admit precise distributional characterization that allows uncertainty quantification on the item parameters, e.g., construction of confidence intervals for the item parameters. The main idea underlying $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ is to randomly pair user-item responses to form item-item comparisons. This is carefully designed to reduce the problem size while retaining statistical independence. We also provide empirical evidence of the efficacy of the two new estimators using both simulated and real data.
@@ 1 Introduction
The item response theory (IRT) (ER, 13) is a framework widely used in psychometrics to model the relationship between individuals’ latent traits (such as ability or personality) and their responses to assessments or questionnaires. It is particularly useful in the development, analysis, and scoring of tests and assessments; see the recent survey in CLLY (25) for a statistical account of IRT.
@@
Among statistical models in IRT, the Rasch model (Ras, 60) is a simple but fundamental one for modeling binary responses. Specifically, for a user $t$ (e.g., test-taker) and an item $i$ (e.g., test problem), the Rasch model assumes that the response of user $t$ to item $i$ is binary and obeys
@@
$\mathbb{P}\left[\text{user $t$ ``loses to'' item $i$}\right]=\frac{e^{\theta_{i}^{\star}}}{e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}}},$
@@
where $\zeta_{t}^{\star},\theta_{i}^{\star}\in\mathbb{R}$ are latent traits of user $t$ and item $i$, respectively. The term “loses to” here refers to negative responses, such as answering an exam question incorrectly, writing a negative review of a product, etc. We also call this response a comparison between user $t$ and item $i$.
@@
In this paper, we focus on estimating the item parameters $\bm{\theta}^{\star}$, which is one of the four main statistical tasks surrounding the Rasch model (or IRT more generally) listed in CLLY (25). Estimating the item parameters is practically important. For instance, in education testing, $\bm{\theta}^{\star}$ could reveal the difficulty of the exam questions, while in product reviews, $\bm{\theta}^{\star}$ could reveal the popularity of the products. Various methods have been proposed for estimating the item parameters $\bm{\theta}^{\star}$, including the joint maximum likelihood estimator (JMLE), the marginal maximum likelihood estimator, the conditional maximum likelihood estimator (CMLE), and the spectral estimator recently proposed in NZ (22, 23). We refer readers to a recent article (Rob, 21) for comparisons between different item parameter estimation methods. However, three main gaps remain in tackling item estimation in the Rasch model:
@@
  • Non-asymptotic guarantee. Apart from the recently proposed spectral estimator (NZ, 22, 23), most theoretical guarantees for the likelihood-based estimators are asymptotic. Since all the estimation procedures are necessarily applied with finite samples, the asymptotic guarantee alone fails to inform practitioners about the performance of different estimators when working with a limited number of samples.
@@
  • Sparse observations. It is not uncommon to encounter situations where each user only responds to a handful of questions. This brings the challenge of incomplete or sparse observations. Many methods, such as CMLE, allow incomplete data (Mol, 95), but most of them lack theoretical support in the sparse regime. While the spectral estimator (NZ, 22, 23) is capable of handling incomplete observations, its theory still requires the observations to be relatively dense. We will elaborate on this point later.
@@
  • Uncertainty quantification. Beyond estimation, uncertainty quantification on the item parameters is central to realizing the full potential of the Rasch model. However, existing results do not address this problem under sparse observations. An exception is the recent work by CLOX (23), which is based on joint estimation and inference on the item parameters $\bm{\theta}^{\star}$ and the user parameters $\bm{\zeta}^{\star}$. Their sampling scheme is more restrictive and requires a relatively dense sampling rate.
In light of these gaps, we raise the following question:
@@ 1.1 Main contributions
The main contribution of our work is the proposal of a novel estimator named random pairing maximum likelihood estimator ($\mathrm{RP\text{-}MLE}$ in short) that achieves the two desiderata listed above.
In essence, $\mathrm{RP\text{-}MLE}$ compiles user-item comparisons to item-item comparisons by randomly pairing responses of the same user to different items. This pairing procedure is carefully designed to extract information of the item parameters while retaining statistical independence. After this compilation step, item parameters $\bm{\theta}^{\star}$ are estimated by the MLE $\widehat{\bm{\theta}}$ given the item-item comparisons.
Even when the observations are extremely sparse, $\mathrm{RP\text{-}MLE}$ achieves the following:
  • Regarding estimation, we show that both $\mathrm{RP\text{-}MLE}$ and its bootstrapped version enjoy optimal finite sample $\ell_{\infty}$ error guarantee. Compared to the conventional $\ell_{2}$ error guarantee, the $\ell_{\infty}$ guarantee, as an entrywise guarantee, is more fine-grained. Consequently, we also show that $\mathrm{RP\text{-}MLE}$ can recover the top-$K$ items with optimal sample complexity.
  • While the optimal $\ell_{\infty}$ error guarantee directly yields optimal $\ell_{2}$ guarantee, such guarantee is only correct in an order-wise sense. We provide a refined finite-sample $\ell_{2}$ error guarantee of $\mathrm{RP\text{-}MLE}$ that is precise even in the leading constant.
  • Supplementing the estimation guarantee, we also build an inferential framework based on $\mathrm{RP\text{-}MLE}$ $\widehat{\bm{\theta}}$. More specifically, we precisely characterize the asymptotic distribution of $\widehat{\bm{\theta}}$. This result facilitates several inferential tasks such as hypothesis testing and construction of confidence regions of $\bm{\theta}^{\star}$.
@@ 1.2 Prior art
The item response theory is a popular statistical framework for modeling response data. It often involves a probabilistic model that links categorical responses to latent traits of both users and items. In the early endeavors, Ras (60) introduced the Rasch model studied herein, and LNB (68) describes a more general framework using parametric models. Popular IRT models include the Rasch model, the two-parameter model (2PL), and the three-parameter logistic model (3PL). As response data widely appears in real life, IRT finds application in numerous fields including educational assessment (DC, 10), psychometrics (LNB, 68), political science (VHSA, 20), and medical assessment (FBC, 05). See CLLY (25) for an overview of IRT.
@@
An important statistical question in the Rasch model is to estimate the item parameters. As the Rasch model is an explicit probabilistic model, many estimation methods are based on the principle of maximizing likelihood. For instance, marginal MLE assumes a prior on the user parameters that is either given or optimized within a parametric distribution family. The item parameter is then estimated by maximizing the marginal likelihood. A drawback is that MMLE relies on a good prior. On the other hand, joint MLE (JMLE) makes no distributional assumption and maximizes the joint likelihood w.r.t. both the item and user parameters. However, it is not consistent for estimating the item parameters when the number of items is fixed (Gho, 95). Interested readers may also consult Lin (99) for an overview of other classical estimators.
Several methods are more relevant to our proposed estimator $\mathrm{RP\text{-}MLE}$ as they follow a similar philosophy to form item-item comparisons from user-item responses. Conditional MLE (CMLE) considers the tuple of all items that are related to a user instead of examining the induced item pairs. It maximizes the likelihood condition on the total number of positive response a user has. Theoretically, And (73) has shown that CMLE is asymptotically normal and consistent. However, no non-asymptotic rate in the setting of sparse observation has been established. Pseudo MLE (PMLE) (Zwi, 95) maximizes the sum of the log-likelihood of all pairs of responses from the same users to different items. However, due to the dependence, no satisfying finite sample performance guarantee has been established. Another related approach is the spectral method, in which a Markov chain on the space of items is formed and the item parameters are estimated via the stationary distribution of the Markov chain. The most recent works in this category are NZ (22, 23), which essentially use the same idea as pseudo MLE in forming item-item comparisons.
@@
An informed reader may realize that the Rasch model resembles the Bradley-Terry-Luce (BTL) model (Luc, 59; BT, 52) in the ranking literature. Indeed, one can view the Rasch model as a special case of the BTL model that distinguishes the two groups of users and items, and only allows inter-group comparisons. There has been a recent surge in interest in studying top-$K$ ranking in the BTL model (SY, 99; YYX, 12; CS, 15; JKSO, 16; CFMW, 19; HYTC, 20; CGZ, 22; GSZ, 23; LFL, 23) and its extensions (HX, 23; HXC, 23; FHY24a, ; FHY24b, ; FLWY25b, ; FLWY25a, ), especially under sparse observations of the pairwise responses. Most notably, under a uniform sampling scheme, CFMW (19) shows that (regularized) MLE and spectral methods are both optimal in top-$K$ ranking and GSZ (23) provides inference results for both methods.
Another line of research focuses on non-uniform sampling. HOX (14) and SBB+ (16) each studies the $\ell_{2}$ error of MLE with a general sampling graph. They obtain high probability upper bound and minimax lower bound that match under in some scenarios. In particular, SBB+ (16) extensively discusses the implication of their result in different graph topology. More recently, general sampling graph has also been studied in top-$K$ ranking. Several articles (HX, 23; Che, 23; LSR, 22) investigate the performance of MLE in the BTL model with a general comparison graph and later YCOM (24) improves the analysis to show the optimality of (weighted) MLE for the BTL model in both uniform and semi-random sampling.
@@
For a positive integer $n$, we denote $n]=\{1,2,\ldots,n\}$. For any $\in\mathbb{R}, aba\wedge b$ means the minimum of $a,b$ and $a\vee b$ means the maximum of $a,b$. We use $\overset{\mathrm{a.s}}{\rightarrow}$, $\overset{\mathrm{p}}{\rightarrow}$, and $\overset{\mathrm{d}}{\rightarrow}$ to denote convergence almost surely, in probability, and in distribution respectively. For a symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$, we use $\lambda_{1}(\bm{A})\geq\lambda_{2}(\bm{A})\geq\cdots\geq\lambda_{n}(\bm{A})$ to denote its eigenvalues and $\bm{A}^{\dagger}$ to denote its Moore-Penrose pseudo-inverse. For symmetric matrices $\bm{A},\bm{B}\in\mathbb{R}^{n\times n}$, $\bm{A}\preceq\bm{B}$ means $\bm{B}-\bm{A}$ is positive semidefinite, i.e., $\bm{v}^{\top}(\bm{B}-\bm{A})\bm{v}\geq 0$ for any $\bm{v}\in\mathbb{R}^{n}$. We use $\bm{e}_{i}$ to denote the standard unit vector with $1$ at $i$-th coordinate and 0 elsewhere. Unless specified otherwise, $\log(\cdot)$ denotes the natural log.
2 Problem setup and new estimators
In this section, we first introduce the formal setup of the item parameter estimation problem in the Rasch model. Then we present the news estimator $\mathrm{RP\text{-}MLE}$ and $\mathrm{MRP\text{-}MLE}$ along with the rationale behind its development.
@@ 2.1 Problem setup
To model sparse user-item responses, we assume that $\mathbb{P}[(t,i)\text{ is compared}]=p$ independently for every $(t,i)\in[n]\times[m]$. To put it in the language of graph theory, we denote the associated bipartite comparison graph to be $\mathcal{G}_{X}=(\mathcal{V}_{X},\mathcal{E}_{X})$, where $\mathcal{V}_{X}$ consists of $n$ users and $m$ items. Then essentially, we are assuming that the bipartite graph follows an Erdős-Rényi random model. [^1] Alternatively, we can assume each user responds to $mp$ items uniformly at random. Our estimator and performance guarantee continue to work in this sampling scheme.
Before moving on, we define condition numbers to characterize the range of the latent traits. Let $\kappa_{1}$, $\kappa_{2}$, and $\kappa$ be defined by $\log\;(\kappa_{1})\coloneqq\max_{ij}\{|\theta_{i}^{\star}-\theta_{j}^{\star}|\}$, $\log(\kappa_{2})\coloneqq\max_{ti}\{|\zeta_{t}^{\star}-\theta_{i}^{\star}|\}$, and $\kappa\coloneqq\max\{\kappa_{1},\kappa_{2}\}$, respectively.
@@ 2.2 Random pairing maximum likelihood estimator
Algorithm 1 Random Pairing Maximum Likelihood Estimator ($\mathrm{RP\text{-}MLE}$)
@@
      Randomly split the $m_{t}$ problems taken by tester $t$ into $\lfloor m_{t}/2\rfloor$ pairs of problems.
@@
    Let $\mathcal{E}_{Y}$ be a set of edges defined by $\mathcal{E}_{Y}\coloneqq\{(i,j):\sum_{t=1}^{n}L_{ij}^{t}\geq 1\}$ and let $\mathcal{G}_{Y}=([m],\mathcal{E}_{Y})$. For each $(i,j)\in\mathcal{E}_{Y}$, let $L_{ij}\coloneqq\sum_{t=1}^{n}L_{ij}^{t}$ and $Y_{ij}\coloneqq(1/L_{ij})\sum_{\{t:L_{ij}^{t}=1\}}Y_{ij}^{t}$.
@@
    Compute MLE on $Y_{ij}$, i.e., $\widehat{\bm{\theta}}\coloneqq\arg\min_{\bm{\theta}:\bm{1}_{m}^{\top}\bm{\theta}=0}\mathcal{L}(\bm{\theta})$, where
@@
In this section, we present our main method $\mathrm{RP\text{-}MLE}$; see Algorithm 1. The algorithm can be divided into two parts. The first part—Steps 1 and 2—uses random pairing to compile the observed user-item responses $\bm{X}\in\mathbb{R}^{n\times m}$ to item-item comparisons $\bm{Y}\in\mathbb{R}^{m\times m}$. The second part—Steps 3 and 4—computes a standard MLE on the item-item comparisons. Some intuition regarding the development of $\mathrm{RP\text{-}MLE}$ is in order.
@@
The idea of pairing is that by matching the responses $X_{ti}$ with $X_{tj}$, we form a comparison between items $i$ and $j$ to directly extract information of item parameters $\theta_{i}^{\star}$ and $\theta_{j}^{\star}$. More specifically, the item-item comparisons $\bm{Y}$ follow the Bradley-Terry-Luce model (BT, 52; Luc, 59), i.e., $\mathbb{P}[Y_{ij}^{t}=1]=e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})$. For the claims made in this part, please see Section A.1 for a formal argument.
@@
More importantly, the pairing is performed in a disjoint fashion. This ensures that all constructed item-item comparisons $Y_{ij}^{t}$ are independent with each other; see Section A.1 for a formal statement. This is the key ingredient that enables us to improve over previous implementation of item-item comparisons, such as pseudo-likelihood (Cho, 82; Zwi, 95) and spectral methods (NZ, 22, 23).
@@
A drawback of this random pairing is that it potentially induces a loss of information since not every possible pairing is considered. Once $X_{ti}$ is paired with $X_{tj}$, we cannot pair $X_{ti}$ with another response $X_{tl}$. That being said, we will later show that the $\ell_{\infty}$ error of $\mathrm{RP\text{-}MLE}$ is still rate-optimal up to logarithmic factors. Hence the loss of information can at most incur a small constant factor in terms of estimation error. Nevertheless, we provide a remedy to this phenomenon in $\mathrm{MRP\text{-}MLE}$ (Algorithm 2) by running (in other words, bootstrapping) the $\mathrm{RP\text{-}MLE}$ multiple times with different random data splitting and averaging the resulting estimates. $\mathrm{MRP\text{-}MLE}$ trivially enjoys the same non-asymptotic $\ell_{\infty}$ error rate (cf. Theorem 1) while improving the estimation error in practice over $\mathrm{RP\text{-}MLE}$. See Figure 4 in Section 4.1 for the empirical evidence.
Algorithm 2 Multiple Random Pairing Maximum Likelihood Estimator ($\mathrm{MRP\text{-}MLE}$)
@@
    Let $n_{\mathrm{split}}$ be the number of runs. For $i=1,\ldots,n_{\mathrm{split}}$, run $\mathrm{RP\text{-}MLE}$ (Algorithm 1), each time with an independent random splitting in Step 1. Let the estimated latent scores be $\widehat{\bm{\theta}}^{(i)}$.
@@
    $\widehat{\bm{\theta}}_{\mathrm{MRP}}=\frac{1}{n_{\mathrm{split}}}\sum_{i=1}^{n_{\mathrm{split}}}\widehat{\bm{\theta}}_{(i)}.$
@@
    Estimate the top-$K$ items by selecting the top-$K$ entries of $\widehat{\bm{\theta}}_{\mathrm{MRP}}$.